Solve each problem. Use your choice of methods.
6x - 6y = -30
24x - 24y = -120

Answer by checkley71(8403)

6X-6Y=-30 & 24X-24Y=-120 MULTIPLY EQUATION #1 BY 6 & SUBTRACT WE GET
24X-24Y=-120
24X-24Y=-120
-----------------
0X 0Y = 0 NO SOLUTION.